artlover03:

Find the surface area of this rectangular prism. Be sure to include the correct unit in your answer.

artlover03:

1 attachment
smackz:

the formula u can use to find the surface area of that rectangular prism is A=2(wl+hl+hw)=2·(4·6+6·6+6·4)= what do you thin kthe answer is?

smackz:

@artlover03

artlover03:

168? i guess im kinda confused @smackz

smackz:

actually you got it right

smackz:

heres the oriogional formula A=2(wl+hl+hw)

smackz:

all i did was take the values of the height , width, and length and plugged them in the formula to get 168 as surface area

smackz:

srry if i confused you

artlover03:

No no! you didnt, im just slow lol

smackz:

it says yds not ft in the pic o;

artlover03:

WAIT I TYPED IT WRONG THE UNITS USED WAS YD... so would it be yd^2? sorry sorry sorry

smackz:

actually it would be yd^3

smackz:

A cube and a square prism are both special types of a rectangular prism. Keep in mind, a square is just a special type of rectangle! Cubes are rectangular prisms where all three dimensions (length, width and height) have the same measurement.

artlover03:

Oh okay, i always get confused. Thank you so much for helping me!

smackz:

np have a nice day :D

artlover03:

Likewise :)